The table shows the demand and supply schedules for magazines. Complete the following sentences. The equilibrium price of a magazine is 4and the equilibrium quantity is 150 magazines a week. Price (dollars per Quantity demanded Quantity supplied 3.00 3.50 4.00 4.50 5.00 160 155 150 145 140 138 144 150 156 161 Now a fall in the price of a newspaper decreases the quantity demanded by 11As the market returns to equilibrium, the quantity demandedand the magazines a week at each price. quantity supplied . At the original equilibrium price, a of a magazine occurs. To return to equilibrium, the price A. shortage; rises B. surplus; rises C. surplus; falls O A. increases: decreases O B. increases; increases O c. decreases; increases O D. decreases; decreases O D. shortage: falls The new equilibrium price is S□ a magazine. Click to select your answer(s).
Answer #1

1.Ans: D) shortage;falls

Explanation:

Original equilibrium price is $4 where quantity demanded equals to quantity supplied (150 units).

Now, due to the change in price of Newspaper the quantity demanded of Magazines decreases by 11 units.

At price $4 , Quantity demanded = 150 - 11 =139

Quantity supplied remains constant.

So there is a shortage.

2.Ans: A) increases; decreases

3.Ans: The new equilibrium price is $3.50 a magazine .

Explanation:

At new equilibrium price $3.50,

Quantity demand =155 - 11 = 144

Quantity supplied = 144 ( it decreases from 150)